OpenStudy (anonymous):

Sam decides to build a square garden. If the area of the garden is 4x2 − 20x + 25 square feet, what is the length of one side of the garden?

OpenStudy (anonymous):

factorize 4x^2 -20x +25.. u will get x=2.5 so length is 2.5

OpenStudy (anonymous):

or if u want u can use the quadratic formula instead of factorizing as factorizing is harder when decimal places are involved

OpenStudy (anonymous):

u know whats the quadratic formula?

OpenStudy (anonymous):

\[x^{} = \frac{ -b + or - \sqrt[]{b^2-4ac} }{ 2a }\]